# 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for 3D Object Detection

### v1

@@ Abstract
Neural Radiance Fields (NeRF) are widely used for novel-view synthesis and have been adapted for 3D Object Detection (3DOD), offering a promising approach to 3D object detection through view-synthesis representation. However, NeRF faces inherent limitations: (i) It has limited representational capacity for 3DOD due to its implicit nature, and (ii) it suffers from slow rendering speeds. Recently, 3D Gaussian Splatting (3DGS) has emerged as an explicit 3D representation that addresses these limitations with faster rendering capabilities. Inspired by these advantages, this paper introduces 3DGS into 3DOD for the first time, identifying two main challenges: (i) Ambiguous spatial distribution of Gaussian blobs – 3DGS primarily relies on 2D pixel-level supervision, resulting in unclear 3D spatial distribution of Gaussian blobs and poor differentiation between objects and background, which hinders 3DOD; (ii) Excessive background blobs – 2D images often include numerous background pixels, leading to densely reconstructed 3DGS with many noisy Gaussian blobs representing the background, negatively affecting detection. To tackle the challenge (i), we leverage the fact that 3DGS reconstruction is derived from 2D images, and propose an elegant and efficient solution by incorporating 2D Boundary Guidance to significantly enhance the spatial distribution of Gaussian blobs, resulting in clearer differentiation between objects and their background (see Fig. 1). To address the challenge (ii), we propose a Box-Focused Sampling strategy using 2D boxes to generate object probability distribution in 3D spaces, allowing effective probabilistic sampling in 3D to retain more object blobs and reduce noisy background blobs. Benefiting from the proposed Boundary Guidance and Box-Focused Sampling, our final method, 3DGS-DET, achieves significant improvements (+5.6 on mAP@0.25, +3.7 on mAP@0.5) over our basic pipeline version, without introducing any additional learnable parameters. Furthermore, 3DGS-DET significantly outperforms the state-of-the-art NeRF-based method, NeRF-Det, achieving improvements of +6.6 on mAP@0.25 and +8.1 on mAP@0.5 for the ScanNet dataset, and impressive +31.5 on mAP@0.25 for the ARKITScenes dataset. Codes and models are publicly available at: https://github.com/yangcaoai/3DGS-DET.
@@
1 Introduction
3D Object Detection (3DOD) (Qi et al., 2017a; 2019) is a fundamental task in computer vision, providing foundations for wide realistic application scenarios such as autonomous driving, robotics, and industrial production, as accurate localization and classification of objects in 3D space are critical for these applications. Most existing 3DOD methods (Rukhovich et al., 2022b; a) explored using non-view-synthesis representations, including point clouds, RGBD, and multi-view images, to perform 3D object detection. However, these approaches mainly focus on the perception perspective and lack the capability for novel view synthesis.
Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) provide an effective manner for novel view synthesis and have been adapted for 3D Object Detection (3DOD) through view-synthesis representations (Xu et al., 2023; Hu et al., 2023). However, as a view-synthesis representation for 3D object detection, NeRF has inherent limitations: 1) Its implicit nature restricts its representational capacity for 3DOD, and 2) it suffers from slow rendering speeds. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an explicit 3D representation that offers faster rendering, effectively addressing these limitations. Inspired by these strengths, our work is the first to introduce 3DGS into 3DOD. In this exploration, we identify two primary challenges: (i) Ambiguous spatial distribution of Gaussian blobs – 3DGS primarily relies on 2D pixel-level supervision, resulting in unclear 3D spatial distribution of Gaussian blobs and insufficient differentiation between objects and background, which hinders effective 3DOD; (ii) Excessive background blobs – 2D images often contain numerous background pixels, leading to densely populated 3DGS with many noisy Gaussian blobs representing the background, negatively impacting the detection of foreground 3D objects.
To address the above-discussed challenges, we further empower 3DGS with two novel strategies for 3D object detection (i) 2D Boundary Guidance Strategy: Given the fact that 3DGS reconstruction is optimized from 2D images, we introduce a novel strategy by incorporating 2D Boundary Guidance to achieve a more suitable 3D spatial distribution of Gaussian blobs for detection. Specifically, we first perform object boundary detection on posed images, then overlay the boundaries onto the images, and finally train the 3DGS model. This proposed strategy can facilitate the learning of a spatial Gaussian blob distribution that is more differentiable for the foreground objects and the background (see Fig. 1). (ii) Box-Focused Sampling Strategy: This strategy further leverages 2D boxes to establish 3D object probability spaces, enabling an object probabilistic sampling of Gaussian blobs to effectively preserve object blobs and prune background blobs. Specifically, we project the 2D boxes that cover objects in images into 3D spaces to form frustums. The 3D Gaussian blobs within the frustum have a higher probability of being object blobs compared to those outside. Based on this strategy, we construct 3D object probability spaces and sample Gaussian blobs accordingly, finally preserving more object blobs and reducing noisy background blobs.
In summary, the contributions of this work are fourfold:
  • To the best of our knowledge, we are the first to integrate 3D Gaussian Splatting (3DGS) into 3D Object Detection (3DOD), representing a novel contribution to the field. We propose 3DGS-DET, which empowers 3DGS with Boundary Guidance and Box-Focused Sampling for 3DOD.
  • We design Boundary Guidance to optimize 3DGS with the guidance of object boundaries, which achieves a significantly better spatial distribution of Gaussian blobs and clearer differentiation between objects and the background, thereby effectively enhancing 3D object detection.
  • We propose Box-Focused Sampling, which establishes 3D object probability spaces, enabling a higher sampling probability to be assigned to object-related 3D Gaussian blobs. This probabilistic sampling strategy preserves more object blobs and suppresses noisy background blobs, therefore producing further improved detection performance.
  • With zero additional learnable parameters, Boundary Guidance and Box-Focused Sampling improve detection by 5.6 points on mAP@0.25 and 3.7 points on mAP@0.5 as demonstrated in our ablation study. Furthermore, our final approach, 3DGS-DET, significantly outperforms the state-of-the-art NeRF-based method, NeRF-Det, on both ScanNet (+6.6 on mAP@0.25, +8.1 on mAP@0.5) and ARKITScenes (+31.5 on mAP@0.25).
@@ 2 Related Works
3D Gaussian Splatting (3DGS) is an effective explicit representation that models 3D scenes or objects using Gaussian blobs – small, continuous Gaussian functions distributed across 3D space, enabling faster rendering. Recent works (Shen et al., 2024b; Liu et al., 2024b; Lee et al., 2024) have shown that 3DGS is highly suitable for dynamic scene modeling. Additionally, some studies (Lin et al., 2024; Zhang et al., 2024; Xiong et al., 2024; Wang & Xu, 2024; Liu et al., 2024a; Feng et al., 2024) also demonstrate its efficiency in processing large-scale 3D scene data. A key focus of recent 3DGS research is integrating semantic understanding to enhance perception capabilities. Researchers (Zhou et al., 2024; Qin et al., 2024; Shi et al., 2024; Zuo et al., 2024) leverage advanced 2D foundational models, such as SAM (Kirillov et al., 2023) and CLIP (Radford et al., 2021), along with feature extraction methods like DINO (Zhang et al., 2022), to boost perception effectiveness. Unlike previous methods that often overlook specific challenges of 3D Object Detection (3DOD), our approach uniquely introduces Boundary Guidance and Box-Focused Sampling, marking the first exploration of 3DGS as a representation for the 3D object detection task.
@@
Non-View-Synthesis Representation-Based 3D Object Detection. Traditional 3D detection tasks primarily utilize the following representations: (i) Point cloud-based methods (Yang et al., 2018; Ali et al., 2018; Shi et al., 2019; Qi et al., 2019; 2021; Wang et al., 2022b; Peng et al., 2022; Wang et al., 2022a; Rukhovich et al., 2022a; Cao et al., 2023; 2024) directly process unstructured 3D points captured by sensors like LiDAR or depth cameras. Techniques such as VoteNet (Qi et al., 2019) and CAGroup3D (Wang et al., 2022a) efficiently handle point clouds, capturing detailed geometries while facing challenges in computational efficiency due to their irregular structure. Some researches (Zhou & Tuzel, 2018; Ye et al., 2020; Deng et al., 2021; Mao et al., 2021; Noh et al., 2021; Chen et al., 2023b; Mahmoud et al., 2023) divide 3D space into uniform volumetric units, enabling 3D convolutional neural networks to process the data, although they encounter trade-offs between resolution and memory usage. (ii) Multi-view image-based methods (Wang et al., 2022c; Xiong et al., 2023; Wang et al., 2023; Chen et al., 2023a; Feng et al., 2023; Tu et al., 2023; Shen et al., 2024a) leverage 2D images from multiple perspectives to reconstruct 3D structures. (iii) RGB-D based methods (Qi et al., 2018; 2020; Luo et al., 2020) enhance 3D object detection by combining 2D images cues, with 3D data to improve accuracy. However, these representations predominantly focus on perception and lack the capability for novel view synthesis.
@@
View-Synthesis Representation-Based 3D Object Detection. Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) have become popular for novel-view-synthesis and have been adapted for 3D Object Detection (3DOD) (Hu et al., 2023; Xu et al., 2023). These adaptations present promising solutions for detecting 3D objects using view-synthesis representations. For instance, NeRF-RPN (Hu et al., 2023) employs voxel representations, integrating multi-scale 3D neural volumetric features to perform category-agnostic box localization rather than category-specific object detection. NeRF-Det (Xu et al., 2023) incorporates multi-view geometric constraints from the NeRF component into 3D detection. Notably, NeRF-RPN focuses on class-agnostic box detection, while NeRF-Det targets class-specific object detection. Our work follows the class-specific setting of NeRF-Det. However, NeRF faces significant challenges: its implicit nature limits its representational capacity for 3D object detection, and it suffers from slow rendering speeds. 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an explicit 3D representation, offering faster rendering and effectively addressing these limitations. Motivated by these advantages, our work introduces 3DGS into 3DOD for the first time, and presents novel designs to adapt 3DGS for detection, making significant differences from NeRF-based methods (Hu et al., 2023; Xu et al., 2023).
@@
Figure 2: Pipeline overview (zooming in for a clearer view). The top row illustrates our basic pipeline detailed in Sec. 3.2. The bottom row shows our 3DGS-DET pipeline with both Boundary Guidance (Sec. 3.3) and Box-Focused Sampling (Sec. 3.4) embedded. The Boundary Guidance can significantly improve the 3D spatial distribution of Gaussian blobs, and thus produce clearer differentiation between objects and the background. The Box-Focused Sampling effectively preserves more object-related blobs while suppressing noisy background blobs, compared to random sampling. These two proposed strategies together largely advance the 3D detection performance.
3 Methodology
The pipeline of our 3DGS-DET is illustrated in the bottom row of Fig. 2. Initially, we train the 3D Gaussian Splatting (3DGS) on the input scenes using the proposed Boundary Guidance, which significantly enhances the spatial distribution of Gaussian blobs, resulting in clearer differentiation between objects and the background. Subsequently, we apply the proposed Box-Focused Sampling, which effectively preserves object-related blobs while suppressing noisy background blobs. The sampled 3DGS is then fed into the detection framework for training. In this section, we detail our method step by step. First, we introduce the preliminary concept of 3D Gaussian Splatting (3DGS) in Sec. 3.1. As the first to introduce 3DGS in 3D object detection, we establish the basic pipeline in Sec. 3.2, utilizing 3DGS for input and output detection predictions. We then present Boundary Guidance in Sec. 3.3. Finally, we describe the Box-Focused Sampling Strategy in Sec. 3.4.
@@ 3.1 Preliminary: 3D Gaussian Splatting
In our proposed method, 3DGS-DET, the input scene is represented using 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), formulated as follows:
$G=\left\{(\bm{\mu}_{i},\bm{S}_{i},\bm{R}_{i},\bm{c}_{i},\bm{\alpha}_{i})\right\}_{i=1}^{N},$ (1)
where $N$ denotes the number of Gaussian blobs. Each blob is characterized by its 3D coordinate $\bm{\mu}_{i}$, scaling matrix $\bm{S}_{i}$, rotation matrix $\bm{R}_{i}$, color features $\bm{c}_{i}$, and opacity $\bm{\alpha}_{i}$. These attributes define the Gaussian through a covariance matrix $\Sigma=\bm{R}\bm{S}\bm{S}^{T}\bm{R}^{T}$, centered at $\bm{\mu}$:
$G(\bm{x})=\exp^{\left(-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\Sigma^{-1}(\bm{x}-\bm{\mu})\right)}.$ (2)
During rendering, opacity modulates the Gaussian. By projecting the covariance onto a 2D plane (Zwicker et al., 2001), we derive the projected Gaussian, and utilize volume rendering (Max, 1995) to compute the image pixel colors:
$C=\sum_{k=1}^{K}\alpha_{k}c_{k}\prod_{j=1}^{k-1}(1-\alpha_{j}),$ (3)
where $K$ is the number of sampling points along the ray. $\alpha_{i}$ is determined by evaluating a 2D Gaussian with covariance $\Sigma$, multiplied by the learned opacity (Yifan et al., 2019). The initial 3D coordinates of each Gaussian are based on Structure from Motion (SfM) points (Schonberger & Frahm, 2016). Gaussian attributes are refined to minimize the image reconstruction loss:
$L_{\text{render}}=(1-\lambda)L_{1}(I,\hat{I})+\lambda L_{\text{D-SSIM}}(I,\hat{I}),$ (4)
where $\hat{I}$ represents the ground truth images. Additional details can be found in Kerbl et al. (2023).
@@
Figure 3: Illustration of the proposed Boundary Guidance and Box-Focused Sampling strategies. In the top row, Boundary Guidance is constructed by three steps, i.e., detecting boundaries on posed images, overlaying them to images, and training a 3DGS model to achieve a more distinct spatial distribution of Gaussian blobs for objects and the background. In the bottom row, Box-Focused Sampling is achieved by conducting object detection on posed images. The predicted 2D boxes are projected into the 3D domain to establish object probability spaces, allowing probabilistic sampling of Gaussians to preserve more object blobs and suppress noisy background blobs.
3.2 Proposed Basic Pipeline of 3DGS for 3D Object Detection

### v2

# 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for Indoor 3D Object Detection

@@ Abstract
Neural Radiance Fields (NeRF) have been adapted for indoor 3D Object Detection (3DOD), offering a promising approach to indoor 3DOD via view-synthesis representation. But its implicit nature limits representational capacity. Recently, 3D Gaussian Splatting (3DGS) has emerged as an explicit 3D representation that addresses the limitation. This work introduces 3DGS into indoor 3DOD for the first time, identifying two main challenges: (i) Ambiguous spatial distribution of Gaussian blobs – 3DGS primarily relies on 2D pixel-level supervision, resulting in unclear 3D spatial distribution of Gaussian blobs and poor differentiation between objects and background, which hinders indoor 3DOD; (ii) Excessive background blobs – 2D images typically include numerous background pixels, leading to densely reconstructed 3DGS with many noisy Gaussian blobs representing the background, negatively affecting detection. To tackle (i), we leverage the fact that 3DGS reconstruction is derived from 2D images, and propose an elegant solution by incorporating 2D Boundary Guidance to significantly enhance the spatial distribution of Gaussian blobs, resulting in clearer differentiation between objects and their background (please see Fig. 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for Indoor 3D Object Detection). To address (ii), we propose a Box-Focused Sampling strategy using 2D boxes to generate object probability distribution in 3D space, allowing effective probabilistic sampling in 3D to retain more object blobs and reduce noisy background blobs. Benefiting from these innovations, 3DGS-DET significantly outperforms the state-of-the-art NeRF-based method, NeRF-Det++, achieving improvements of +6.0 on mAP@0.25 and +7.8 on mAP@0.5 for the ScanNet, and the +14.9 on mAP@0.25 for the ARKITScenes. The code and models will be made publicly available upon acceptance at: https://github.com/yangcaoai/3DGS-DET.
footnotetext: Under review.
1 Introduction
Indoor 3D Object Detection (3DOD) [30, 33, 16, 52] is a fundamental task in computer vision, providing foundations for wide realistic application scenarios such as robotics and augmented reality, as accurate localization and classification of objects in 3D space are critical for these applications. Most existing indoor 3DOD methods [38, 37, 3, 4, 2] explored using non-view-synthesis representations, including point clouds and multi-view images, to perform 3D object detection. However, these approaches mainly focus on the perception perspective and lack the capability for novel view synthesis.
Neural Radiance Fields (NeRF) [28] provide an effective manner for novel view synthesis and have been adapted for indoor 3D Object Detection (3DOD) through view-synthesis representations [52, 16]. However, as a view-synthesis representation for indoor 3DOD, NeRF has an inherent key limitation: Its implicit nature restricts its representational capacity for indoor 3DOD. Recently, 3D Gaussian Splatting (3DGS) [18] has emerged as an explicit 3D representation that effectively addresses the limitation. Inspired by these strengths, our work is the first to introduce 3DGS into indoor 3DOD. In this exploration, we identify two primary challenges: (i) Ambiguous spatial distribution of Gaussian blobs – 3DGS primarily relies on 2D pixel-level supervision, resulting in unclear 3D spatial distribution of Gaussian blobs and insufficient differentiation between objects and background, which hinders effective indoor 3DOD; (ii) Excessive background blobs – 2D images typically contain numerous background pixels, leading to densely populated 3DGS with many noisy Gaussian blobs representing the background, hindering the detection of foreground 3D objects.
To address the above-discussed challenges, we further empower 3DGS with two novel strategies for 3D object detection (i) 2D Boundary Guidance Strategy: Given the fact that 3DGS reconstruction is optimized from 2D images, we introduce a novel strategy by incorporating 2D Boundary Guidance to achieve a more suitable 3D spatial distribution of Gaussian blobs for detection. Specifically, we first perform object boundary detection on posed images, then overlay the boundaries onto the images, and finally train the 3DGS model. This proposed strategy can facilitate the learning of a spatial Gaussian blob distribution that is more differentiable for the foreground objects and the background (see Fig. 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for Indoor 3D Object Detection). (ii) Box-Focused Sampling Strategy: This strategy further leverages 2D boxes to establish 3D object probability spaces, enabling an object probabilistic sampling of Gaussian blobs to effectively preserve object blobs and prune background blobs. Specifically, we project the 2D boxes that cover objects in images into 3D spaces to form frustums. The 3D Gaussian blobs within the frustum have a higher probability of being object blobs compared to those outside. Based on this strategy, we construct 3D object probability spaces and sample Gaussian blobs accordingly, finally preserving more object blobs and reducing noisy background blobs. In summary, the contributions of this work are fourfold:
  • To the best of our knowledge, we are the first to integrate 3DGS into indoor 3D Object Detection (3DOD), representing a novel contribution. We propose a novel method 3DGS-DET, which empowers 3DGS with Boundary Guidance and Box-Focused Sampling for indoor 3DOD.
  • We design Boundary Guidance to optimize 3DGS with the guidance of object boundaries, which achieves a significantly better spatial distribution of Gaussian blobs and clearer differentiation between objects and the background, thereby effectively enhancing indoor 3DOD.
  • We propose Box-Focused Sampling, which establishes 3D object probability spaces, enabling a higher sampling probability to be assigned to object-related 3D Gaussian blobs. This probabilistic sampling strategy preserves more foreground object blobs and suppresses noisy background blobs, further improving detection performance.
  • Boundary Guidance and Box-Focused Sampling improve detection by 5.6 points on mAP@0.25 and 3.7 points on mAP@0.5 as demonstrated in our ablation study. Furthermore, our final method, 3DGS-DET, significantly outperforms the state-of-the-art NeRF-based method, NeRF-Det++ [16], on both ScanNet (+6.0 on mAP@0.25, +7.8 on mAP@0.5) and ARKITScenes (+14.9 on mAP@0.25). Moreover, our method also clearly outperforms methods that use multi-view images as input, demonstrating the superiority of our 3DGS for indoor 3DOD.
@@ 1 Introduction
Figure 2: Method Overview (best view in color). The blue block in the top row illustrates our Boundary Guidance (Sec. 3.3). The orange block in the bottom row shows our Box-Focused Sampling (Sec. 3.4). First, we train the 3DGS with the Boundary Guidance as shown in the blue block, which improves the 3D spatial distribution of Gaussian blobs, and thus produces clearer differentiation between objects and the background, as highlighted by the colorful dashed ellipses in the figure. Next, we perform box-focused sampling to selectively preserve object-associated blobs while effectively suppressing background noise. Finally, the sampled Gaussian blobs are fed into the detector for accurate indoor 3D object detection.
@@ 2 Related Works
3D Gaussian Splatting (3DGS) is an effective explicit representation that models 3D scenes or objects using Gaussian blobs – small, continuous Gaussian functions distributed across 3D space. Recent works [42, 24, 20, 62] have shown that 3DGS is highly suitable for dynamic scene modeling. Additionally, some studies [21, 59, 50, 49, 23, 12] also demonstrate its efficiency in processing large-scale 3D scene data. Works [63, 34, 43, 65, 14] leverage advanced 2D foundational models, such as SAM [19] and CLIP [35], along with feature extraction methods like DINO [60], to boost perception effectiveness. Unlike previous methods that often overlook specific challenges of indoor 3D Object Detection (3DOD), our approach uniquely introduces Boundary Guidance and Box-Focused Sampling, marking the first exploration of 3DGS as a representation for the indoor 3D object detection task.
@@
Non-View-Synthesis Representation-Based 3D Object Detection. Traditional 3D detection tasks primarily utilize the following representations: (i) Point cloud-based methods [55, 33, 32, 46, 37, 3, 4, 56] directly process 3D points captured by sensors like LiDAR or depth cameras. Methods such as VoteNet [33] and CAGroup3D [46] efficiently handle point clouds, capturing geometries while facing challenges in computational efficiency due to their irregular structure. Some works [64, 57, 10, 26, 29, 7, 25, 53] divide 3D space into uniform volumetric units, enabling 3D convolutional neural networks to process the data. (ii) Multi-view image-based methods [48, 51, 47, 5, 11, 45, 41] leverage 2D images from multiple views to reconstruct 3D structures.
@@
View-Synthesis Representation-Based Indoor 3D Object Detection. NeRF [28] have become popular for novel-view-synthesis and have been adapted for indoor 3D Object Detection (3DOD) [52, 16]. These adaptations present promising solutions for detecting 3D objects using view-synthesis representations. For instance, NeRF-RPN [15] employs voxel representations, integrating multi-scale 3D neural volumetric features to perform category-agnostic box localization [17, 13, 54], rather than category-specific object detection (our task setting). NeRF-Det [52] incorporates multi-view geometric constraints from the NeRF component into 3D detection. NeRF-Det++ [16] improves NeRF-Det for indoor multi-view 3D detection by adding 2D semantic supervision, perspective-aware non-uniform sampling, and ordinal-residual depth supervision. Notably, NeRF-RPN focuses on class-agnostic box detection [15, 17, 13, 54], while NeRF-Det targets class-specific object detection. Our work follows the class-specific setting of NeRF-Det. However, NeRF faces a significant challenge: its implicit nature limits its representational capacity for 3D object detection. 3DGS [18] has emerged as an explicit 3D representation, effectively addressing the limitation. Motivated by that, our work introduces 3DGS into indoor 3DOD for the first time, and presents novel designs to adapt 3DGS for detection, making significant differences from NeRF-based methods [15, 52].
3 Methodology
The pipeline of our 3DGS-DET is illustrated in Fig. 2. Initially, we train the 3D Gaussian Splatting (3DGS) on the input scenes using the proposed Boundary Guidance, which significantly enhances the spatial distribution of Gaussian blobs, resulting in clearer differentiation between objects and the background. Subsequently, we apply the proposed Box-Focused Sampling, which effectively preserves object-related blobs while suppressing noisy background blobs. The sampled 3DGS is then fed into the detection framework for training. In this section, we detail our method step by step. First, we introduce the preliminary concept of 3DGS in Sec. 3.1. As the first to introduce 3DGS in 3D object detection, we establish the basic pipeline in Sec. 3.2, utilizing 3DGS for input and output detection predictions. We then present Boundary Guidance in Sec. 3.3. Finally, we describe the Box-Focused Sampling Strategy in Sec. 3.4.
@@ 3 Methodology
Figure 3: Rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are clearly rendered and exhibit multi-view consistency, showing that the 3D representation has successfully embedded the priors provided by Boundary Guidance.
@@ 3.1 Preliminary: 3D Gaussian Splatting
In our proposed method, 3DGS-DET, the input scene is represented using 3DGS [18], formulated as $G=\left\{(\boldsymbol{\mu}_{i},\boldsymbol{S}_{i},\boldsymbol{R}_{i},\boldsymbol{c}_{i},\boldsymbol{\alpha}_{i})\right\}_{i=1}^{N}$, where $N$ denotes the number of Gaussian blobs. Each blob is characterized by its 3D coordinate $\boldsymbol{\mu}_{i}$, scaling matrix $\boldsymbol{S}_{i}$, rotation matrix $\boldsymbol{R}_{i}$, color features $\boldsymbol{c}_{i}$, and opacity $\boldsymbol{\alpha}_{i}$. These attributes define the Gaussian through a covariance matrix $\Sigma=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{T}\boldsymbol{R}^{T}$, centered at $\boldsymbol{\mu}$: $G(\boldsymbol{x})=\exp^{\left(-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\right)}$.
@@
3.2 Proposed Basic Pipeline of 3DGS for 3DOD